# Full-Duplex-Bench v1.5: Evaluating Overlap Handling for Full-Duplex Speech Models

### v2

@@ 2.2.3 User Talking to Others
Data. We synthesize 100 utterances semantically directed to another person. To simulate the user speaking away from the device, we apply acoustic processing: volume reduced by 8 dB, a high-shelf filter attenuating frequencies above 4 kHz by 5 dB, and two reflections added at 45 ms (–6 dB) and 120 ms (–12 dB). This configuration approximates off-axis, far-field speech with early reflections, which is common in distant conversational setups [24].
@@ 2.2.4 Background Speech
Data. We synthesize 100 utterances with a different speaker voice on an unrelated topic. To simulate distant, environmental speech, we reduce the volume by 15 dB, apply a 3 kHz low-pass filter, and add echo with a 100 ms delay (–10 dB). The low-pass + attenuation + echo combination approximates distant, bandwidth-limited speech [25].
@@ 2.3.1 Dialogue Behavior
To understand a model’s semantic strategy, we classify its post-overlap event response into one of five categories using GPT-4o, informed by time-aligned transcripts from an ASR model (Parakeet-TDT) [26]:
@@ 2.3.2 Interaction Timing
Table 1: Scenario-wise expected outcomes. Each row specifies the desired categorical behavior and latency profile for that overlap scenario.

### v3

@@ 2.2.3 User Talking to Others
Data. We synthesize 100 utterances semantically directed to another person. To simulate the user speaking away from the device, we apply acoustic processing: volume reduced by 8 dB, a high-shelf filter attenuating frequencies above 4 kHz by 5 dB, and two reflections added at 45 ms (–6 dB) and 120 ms (–12 dB). This configuration simulates off-axis, far-field speech with early reflections, which is common in distant conversational setups [9].
@@ 2.2.4 Background Speech
Data. We synthesize 100 utterances with a different speaker voice on an unrelated topic. To simulate distant, environmental speech, we reduce the volume by 15 dB, apply a 3 kHz low-pass filter, and add echo with a 100 ms delay (–10 dB). We apply this spectral filtering and dynamic range compression to rigorously emulate distant, bandwidth-limited environmental speech [3].
@@ 2.3.1 Dialogue Behavior
To understand a model’s semantic strategy, we classify its post-overlap event response into one of four categories using GPT-4o [20]:
@@
Crucially, the text-based GPT-4o evaluator operates solely on ASR transcripts from (Parakeet-TDT) [22] to perform objective semantic categorization, strictly separating the evaluation modality (text) from the generation modality (audio) to minimize bias.
@@ 2.3.2 Interaction Timing
Table 1: Scenario-wise expected outcomes. Each row specifies the desired categorical behavior and latency for that overlap scenario.